# Deep learning joint extremes of metocean variables using the SPAR model

### v1

@@ Abstract
This paper presents a novel deep learning framework for estimating multivariate joint extremes of metocean variables, based on the Semi-Parametric Angular-Radial (SPAR) model. When considered in polar coordinates, the problem of modelling multivariate extremes is transformed to one of modelling an angular density, and the tail of a univariate radial variable conditioned on angle. In the SPAR approach, the tail of the radial variable is modelled using a generalised Pareto (GP) distribution, providing a natural extension of univariate extreme value theory to the multivariate setting. In this work, we show how the method can be applied in higher dimensions, using a case study for five metocean variables: wind speed, wind direction, wave height, wave period and wave direction. The angular variable is modelled empirically, while the parameters of the GP model are approximated using fully-connected deep neural networks. Our data-driven approach provides great flexibility in the dependence structures that can be represented, together with computationally efficient routines for training the model. Furthermore, the application of the method requires fewer assumptions about the underlying distribution(s) compared to existing approaches, and an asymptotically justified means for extrapolating outside the range of observations. Using various diagnostic plots, we show that the fitted models provide a good description of the joint extremes of the metocean variables considered.
@@ 1 Introduction
Many problems in offshore and coastal engineering require estimation of joint extremes for metocean variables. Responses of offshore and coastal structures are dependent on multiple variables, such as wind speed and direction, wave height, period and direction, current speed and direction. Providing accurate and reliable estimates of the joint extremes in this setting is a challenging problem for metocean engineers. Various design standards recommend the use of the environmental contour method [haver2008environmental]. Some types of contour can be estimated without an explicit model for the joint distribution of variables [Derbanne2019, mackay2023-DIFORM]. However, environmental contour methods typically make simplifying assumptions and only give approximate estimates of long-term extreme responses [haselsteiner2022_NREL, speers2024]. Full probabilistic analysis of long-term extreme responses requires a model for the joint density of the relevant metocean variables. A wide range of approaches have been proposed for estimating joint densities. In the offshore engineering literature, the two most popular approaches are global hierarchical models and copula models – see, e.g. [Ross2020, haselsteiner2021_benchmark].
Let $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$ denote a continuous random vector with joint density function $f_{\mathbf{X}}$, and marginal density and distribution functions $f_{X_{j}}$ and $F_{X_{j}},$ respectively, for $j=1,...,d$. In the global hierarchical approach [bitner2015joint, horn2018new, cheng2019long], the joint density is written as
$f_{\mathbf{X}}(\mathbf{x})=f_{X_{1}}(x_{1})\,f_{X_{2}|X_{1}}(x_{2}|x_{1})\cdots f_{X_{d}|(X_{1},...,X_{d-1})}(x_{d}|(x_{1},...,x_{d-1})),$ (1)
where $f_{X_{j}|(X_{1},...,X_{j-1})}$ is the density of $X_{j}$ conditional on $(X_{1},...,X_{j-1})$ for $j\in\{2,\ldots,d\}$. Inference typically involves selecting parametric forms for $f_{X_{1}}(x_{1})$, $f_{X_{2}|X_{1}}$, …, $f_{X_{d}|(X_{1},...,X_{d-1})}$ and estimating relations between the parameters of the conditional densities and the conditioning variables. There are various problems with this approach. Firstly, there is no a priori reason to suppose that variables follow any particular parametric distribution, and misspecified models can have dramatic consequences when approximating dependence structures. Secondly, a model fit to all of the observations does not guarantee a good fit to the tail, which is the region of interest for extremes. Finally, the models for the parameters of the conditional densities are usually based on ad hoc assumptions, and provide no rationale for extrapolating outside the range of observations. In many cases, it has been shown that such models provide a poor fit to observed data [deHauteclocque2022].
@@
There are also a wide range of methods in the statistical literature for modelling joint extremes (e.g. [Beirlant2004, Ledford1997, Wadsworth2013]). However, many of these approaches make strong assumptions about the dependence structure, or copula, which are often not supported by environmental datasets [huser2024modeling]. The most popular choice for metocean variables is the conditional extremes model [Heffernan2004], which describes the joint distribution of variables conditional on at least one variable being large. The key limitation of this approach is that it only characterises the region of variable space where the conditioning variable is large, and inferences made using different conditioning variables are not necessarily consistent [Liu2014]. A further limitation of this method (and other methods in the multivariate extremes literature) is that it requires a transformation of the margins to a standard scale. This requires first estimating the marginal distributions for each variable – a process which is subject to uncertainty. Furthermore, it has been demonstrated that poor marginal estimates greatly affect the quality of the resulting multivariate inference [towe2023estimation].
@@
In this paper, we discuss the application of a new method, introduced in [mackay2022imex], which overcomes the limitations of existing approaches and provides a general, flexible framework for modelling multivariate extremes. The model is referred to as the Semi-Parametric Angular-Radial (SPAR) model. The SPAR model provides a framework for estimating multivariate extremes that does not require strong assumptions about the form of the margins or dependence structure, and provides a justified means of extrapolating outside the range of observations. Moreover, the model is only fitted to extreme observations, meaning that no assumptions are required about the bulk of the distribution. Theoretical aspects of the SPAR model are presented in [mackay2023-SPAR], and an inference approach in a two-dimensional setting is provided in [murphy2024inference, mackay2025spar]. The purpose of this paper is to extend the modelling method to the general multivariate setting.
@@
The paper is organised as follows. Section 2 describes a brief overview of the theoretical aspects of the model. Our deep learning approach for estimating SPAR model parameters is introduced in Section 3. Section 4 presents an example application of the model to a five-dimensional problem: estimating the joint extremes of wind speed, wave direction, wave height, wave period and wave direction. We discuss the challenges that arise for these particular variables, and how well the model assumptions are satisfied in this setting. We conclude in Section 5 with a discussion and outlook on future work.
@@ 2 Theory
The SPAR model is an extension of the univariate peaks-over-threshold (POT) method to the multivariate setting. It involves a transformation of variables to angular-radial coordinates, and then models the upper-tail of the radial variable, conditional on angle, using a non-stationary generalised Pareto (GP) model. Suppose that we have a continuous random vector $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$ with joint density function $f_{\mathbf{X}}$. We define radial and angular variables as
@@
$f_{R,\mathbf{W}}(r,\mathbf{w})=r^{d-1}\,f_{\mathbf{X}}(r\mathbf{w}),$ (4)
@@
Noting that $\mathbf{X}=R\mathbf{W}$, and that $\mathbf{W}$ lies on the surface of the unit hypersphere, we can see that the ‘extreme’ parts of the distribution of $\mathbf{X}$ correspond to large values of the radial variable at any given angle. Therefore, the problem of modelling multivariate extremes is transformed to that of modelling an angular density $f_{\mathbf{W}}$ and the tail of the conditional radial density $f_{R|\mathbf{W}}$. For a given angle $\mathbf{w}$, the density $f_{R|\mathbf{W}}(r|\mathbf{w})$ is univariate. Univariate extreme value theory suggests that a suitable model for the tail of $f_{R|\mathbf{W}}$ is the GP distribution, with parameters conditional on angle (e.g., [Coles2001]). This motivates the SPAR model, whereby parametric and non-parametric models are used to model the conditional radial and angular distributions, respectively. Define a threshold function $u(\mathbf{w})>0$ to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ at exceedance probability $\zeta\in(0,1)$, with $\zeta$ close to 0, i.e., the solution of $\zeta=\Pr(R>u(\mathbf{w})|\mathbf{W}=\mathbf{w})$. The SPAR model can be written as
$f_{R,\mathbf{W}}(r,\mathbf{w})=\zeta f_{\mathbf{W}}(\mathbf{w})f_{GP}(r-u(\mathbf{w});\xi(\mathbf{w}),\sigma(\mathbf{w})),\quad r>u(\mathbf{w}),$ (6)
where $f_{GP}$ is the GP density function, and $\xi(\mathbf{w})\in\mathbb{R}$ and $\sigma(\mathbf{w})>0$ are shape and scale parameters, respectively, given as functions of the angle $\mathbf{w}$. The GP density function is given by
$f_{GP}(r;\xi,\sigma)=\begin{cases}\dfrac{1}{\sigma}\left(1+\xi\dfrac{r}{\sigma}\right)^{-1-\tfrac{1}{\xi}},&\xi\neq 0,\\[10.0pt] \dfrac{1}{\sigma}\exp\left(-\dfrac{r}{\sigma}\right),&\xi=0,\end{cases}$ (7)
@@
Many non-parametric methods for estimation of densities assume that the density is finite and continuous. Similarly, many representations for non-stationary modelling of parametric distributions assume that the parameter functions are finite and continuous. Therefore, to simplify our inference, we assume that the angular density $f_{\mathbf{W}}$, threshold function $u(\mathbf{w})$, and GP parameter functions, $\xi(\mathbf{w})$ and $\sigma(\mathbf{w})$ are finite and continuous with respect to the angle $\mathbf{w}$.
@@
$f_{\mathbf{X}}(r\mathbf{w})=\zeta\,r^{1-d}\,f_{\mathbf{W}}(\mathbf{w})\,f_{GP}(r-u(\mathbf{w});\xi(\mathbf{w}),\sigma(\mathbf{w})).$ (8)
Calculating marginal and joint probabilities using the SPAR model then involves either integration of the joint density over specified angular and radial domains, or via Monte Carlo techniques, i.e., by simulating from the estimated model and deriving probability estimates empirically. To simulate from the SPAR model, we first draw an angle $\mathbf{w}$ from $f_{\mathbf{W}}$, then use inversion sampling to generate a corresponding radial value $z$ from the GP distribution with parameter vector $(\xi(\mathbf{w}),\sigma(\mathbf{w}))$, and finally define $r=u(\mathbf{w})+z$. The pair $(r,\mathbf{w})$ is then a random sample from the SPAR model. This can be converted back to the original variable space using the inverse transformation $\mathbf{x}=r\mathbf{w}$. As the SPAR model is only fitted to observations for which $r>u(\mathbf{w})$, one can create a sample (of the original random vector $\mathbf{X}$) of size $N$ by simulating $\zeta N$ points from the SPAR model, and then resampling $(1-\zeta)N$ points from observations with $r<u(\mathbf{w})$. The rationale for this is that there should be a sufficient number of observations within the body of the distribution to obtain a reasonable estimate from resampling.
As described in [mackay2025spar], the SPAR model provides an explicit means for calculating a contour with a specified exceedance probability. However, this contour is defined in terms of the probability of an observation falling anywhere outside the region, or the ‘total exceedance probability’. As such, these contours are more conservative than those defined in terms of marginal exceedance probabilities, such as IFORM contours (or variants thereof), with the conservatism increasing with the number of dimensions [mackay2021marginal]. If the primary interest of the analysis is to estimate environmental contours, then the use of the SPAR model is not necessary. Instead, we recommend the use of the Direct-IFORM method [Derbanne2019, mackay2023-DIFORM], which does not require a model for the joint density or any assumptions about the dependence structure between the variables.
@@ 3 Inference
Inference for the SPAR model involves estimating the angular density, and the GP threshold and parameter functions. These problems are separable: inference for $f_{\mathbf{W}}$ can be conducted independently of that for $(u(\mathbf{w}),\sigma(\mathbf{w}),\xi(\mathbf{w}))$. In this section, we discuss inference for the angular density in Section 3.1, then discuss modelling of the conditional radial variable in Section 3.2. Code for fitting our model is available upon reasonable request.

### v2

@@ Abstract
This paper presents a novel deep learning framework for estimating multivariate joint extremes of metocean variables, based on the Semi-Parametric Angular-Radial (SPAR) model. When considered in polar coordinates, the problem of modelling multivariate extremes is transformed to one of modelling an angular density, and the tail of a univariate radial variable conditioned on angle. In the SPAR approach, the tail of the radial variable is modelled using a generalised Pareto (GP) distribution, providing a natural extension of univariate extreme value theory to the multivariate setting. In this work, we show how the method can be applied in higher dimensions, using a case study for five metocean variables: wind speed, wind direction, wave height, wave period, and wave direction. The angular variable is modelled using a kernel density method, while the parameters of the GP model are approximated using fully-connected deep neural networks. Our approach provides great flexibility in the dependence structures that can be represented, together with computationally efficient routines for training the model. Furthermore, the application of the method requires fewer assumptions about the underlying distribution(s) compared to existing approaches, and an asymptotically justified means for extrapolating outside the range of observations. Using various diagnostic plots, we show that the fitted models provide a good description of the joint extremes of the metocean variables considered.
@@ 1 Introduction
Many problems in offshore and coastal engineering require estimation of joint extremes for metocean variables. Responses of offshore and coastal structures are dependent on multiple variables, such as wind speed and direction, wave height, period and direction, flow speed and direction. Providing reliable estimates of the joint extremes in this setting is a challenging problem for metocean engineers. Various design standards recommend the use of the environmental contour method [haver2008environmental]. Some types of contour can be estimated without an explicit model for the joint distribution of variables [Derbanne2019, mackay2023-DIFORM]. However, environmental contour methods typically make simplifying assumptions and only give approximate estimates of long-term extreme responses [haselsteiner2022_NREL, speers2024]. Full probabilistic analysis of long-term extreme responses requires a model for the joint density of the relevant metocean variables. A wide range of approaches have been proposed for estimating joint densities. In the offshore engineering literature, the two most popular approaches are global hierarchical models and copula models – see, e.g., [Ross2020, haselsteiner2021_benchmark].
Let $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$, $d\geq 2$ denote a continuous random vector with joint density function $f_{\mathbf{X}}$, and marginal density and cumulative distribution functions $f_{X_{j}}$ and $F_{X_{j}},$ respectively, for $j=1,...,d$. In the global hierarchical approach [bitner2015joint, horn2018new, cheng2019long, simao2022multi], the joint density is written as
$f_{\mathbf{X}}(\mathbf{x})=f_{X_{1}}(x_{1})\,f_{X_{2}|X_{1}}(x_{2}|x_{1})\cdots f_{X_{d}|(X_{1},...,X_{d-1})}(x_{d}|x_{1},...,x_{d-1}),$ (1)
where $f_{X_{j}|(X_{1},...,X_{j-1})}$ is the density of $X_{j}$ conditional on $(X_{1},...,X_{j-1})$ for $j\in\{2,\ldots,d\}$. Inference typically involves selecting parametric forms for $f_{X_{1}}$, $f_{X_{2}|X_{1}}$, …, $f_{X_{d}|(X_{1},...,X_{d-1})}$ and estimating relations between the parameters of the conditional densities and the conditioning variables. There are various problems with this approach. Firstly, there is no a priori reason to suppose that variables follow any particular parametric distribution, and misspecified models can have dramatic consequences when approximating dependence structures. Secondly, a model fit to all of the observations does not guarantee a good fit to the tail, which is the region of interest for extremes. Finally, the models for the parameters of the conditional densities are usually based on ad hoc assumptions, and provide no rationale for extrapolating outside the range of observations. In many cases, it has been shown that such models provide a poor fit to observed data, especially in extreme regions [deHauteclocque2022].
@@
There are also a wide range of methods in the statistical literature for modelling joint extremes (e.g., [Beirlant2004, Ledford1997, Wadsworth2013]). However, many of these approaches make strong assumptions about the dependence structure, or copula, which are often not supported by environmental datasets [huser2024modeling]. The most popular choice for metocean variables is the conditional extremes model [Heffernan2004], which describes the joint distribution of variables conditional on at least one variable being large. The key limitation of this approach is that it only characterises the region of variable space where the conditioning variable is large, and inferences made using different conditioning variables are not necessarily consistent [Liu2014]. A further limitation of this method (and other methods in the multivariate extremes literature) is that it requires a transformation of the margins to a standard scale. This requires first estimating the marginal distributions for each variable – a process which is subject to uncertainty. Furthermore, it has been demonstrated that poor marginal estimates greatly affect the quality of the resulting multivariate inference [towe2023estimation].
@@
In this paper, we discuss the application of a new method, introduced in [mackay2022imex], which overcomes the limitations of existing approaches and provides a general, flexible framework for modelling multivariate extremes. The model is referred to as the Semi-Parametric Angular-Radial (SPAR) model. The SPAR model provides a framework for estimating multivariate extremes that does not require strong assumptions about the form of the margins or dependence structure, and provides a justified means of extrapolating outside the range of observations. Moreover, the model is only fitted to extreme observations, meaning that no assumptions are required about the bulk of the distribution. Theoretical aspects of the SPAR model are presented in [mackay2023-SPAR], and an inference approach in a two-dimensional setting is provided in [murphy2024inference, mackay2025spar]. The purpose of this paper is to extend the modelling method to the general multivariate setting, with two or more dimensions. The two-dimensional inference scheme in [murphy2024inference, mackay2025spar] utilised cubic regression splines to model the angular dependence of the radial distribution parameters. In this work we take a different approach, and adopt a deep learning scheme, where the radial distribution parameters are modelled using artificial neural networks. As discussed further in Section 3, this offers some computational advantages in higher-dimensional settings.
@@
The paper is organised as follows. Section 2 describes a brief overview of the theoretical aspects of the model. Our deep learning approach for estimating the radial component of the SPAR model is introduced in Section 3. Section 4 presents an example application of the model to a five-dimensional problem: estimating the joint extremes of wind speed, wave direction, wave height, wave period, and wave direction. We discuss the challenges that arise for these particular variables, and how well the model assumptions are satisfied in this setting. Several novel types of diagnostic plots are introduced to assess the fit of the model. We conclude in Section 5 with a discussion and outlook on future work.
@@ 2 Theory
2.1 SPAR model definition
The SPAR model can be viewed as an extension of the univariate peaks-over-threshold (POT) method to the multivariate setting. It involves a transformation of variables to angular-radial coordinates, and then models the upper-tail of the radial variable, conditional on angle, using a non-stationary generalised Pareto (GP) model. Suppose that we have a continuous random vector $\mathbf{X}=(X_{1},...,X_{d})\in\mathbb{R}^{d}$ with joint density function $f_{\mathbf{X}}$. We define radial and angular variables as
@@ 2.1 SPAR model definition
$f_{R,\mathbf{W}}(r,\mathbf{w})=r^{d-1}\,f_{\mathbf{X}}(r\mathbf{w}),\quad r\geq 0,\,\mathbf{w}\in\mathbb{S}^{d-1},$ (4)
@@
Noting that $\mathbf{X}=R\mathbf{W}$, and that $\mathbf{W}$ lies on the surface of the unit hypersphere, we can see that the ‘extreme’ parts of the distribution of $\mathbf{X}$ correspond to large values of the radial variable at any given angle. Therefore, the problem of modelling multivariate extremes is transformed to that of modelling an angular density $f_{\mathbf{W}}$ and the tail of the conditional radial density $f_{R|\mathbf{W}}$. For a given angle $\mathbf{w}$, the density $f_{R|\mathbf{W}}(r|\mathbf{w})$ is univariate. Univariate extreme value theory suggests that a suitable model for the tail of $f_{R|\mathbf{W}}$ is the GP distribution, with parameters conditional on angle (e.g., [Coles2001]). This motivates the SPAR model, whereby parametric and non-parametric models are used to model the conditional radial and angular distributions, respectively. Define a threshold function $u(\mathbf{w})>0$ to be the quantile of $R|(\mathbf{W}=\mathbf{w})$ with exceedance probability $\zeta\in(0,1)$, with $\zeta$ close to 0, i.e., the solution of $\zeta=\Pr(R>u(\mathbf{w})|\mathbf{W}=\mathbf{w})$. The SPAR model can be written as
$f_{R,\mathbf{W}}(r,\mathbf{w})=\zeta f_{\mathbf{W}}(\mathbf{w})f_{\rm GP}(r-u(\mathbf{w});\xi(\mathbf{w}),\sigma(\mathbf{w})),\quad r>u(\mathbf{w}),\,\mathbf{w}\in\mathbb{S}^{d-1},$ (6)
where $f_{\rm GP}$ is the GP density function, and $\xi(\mathbf{w})\in\mathbb{R}$ and $\sigma(\mathbf{w})>0$ are shape and scale parameters, respectively, given as functions of the angle $\mathbf{w}$. The GP density function is given by
$f_{\rm GP}(r;\xi,\sigma)=\begin{cases}\dfrac{1}{\sigma}\left(1+\xi\dfrac{r}{\sigma}\right)^{-1-\tfrac{1}{\xi}},&\xi\neq 0,\\[10.0pt] \dfrac{1}{\sigma}\exp\left(-\dfrac{r}{\sigma}\right),&\xi=0,\end{cases}$ (7)
@@
Many non-parametric methods for estimation of densities, such as kernel density methods, mixture models or spline-based methods, assume that the density is finite and continuous. Similarly, many representations for non-stationary modelling of parametric distributions assume that the parameter functions are finite and continuous. Therefore, to simplify our inference, we assume that the angular density $f_{\mathbf{W}}$, threshold function $u(\mathbf{w})$, and GP parameter functions, $\xi(\mathbf{w})$ and $\sigma(\mathbf{w}),$ are finite and continuous with respect to the angle $\mathbf{w}$.
@@
$f_{\mathbf{X}}(r\mathbf{w})=\zeta\,r^{1-d}\,f_{\mathbf{W}}(\mathbf{w})\,f_{\rm GP}(r-u(\mathbf{w});\xi(\mathbf{w}),\sigma(\mathbf{w})).$ (8)
Calculating marginal and joint probabilities using the SPAR model then involves either integration of the joint density over specified angular and radial domains, or via Monte Carlo techniques, i.e., by simulating from the estimated model and deriving probability estimates empirically. To simulate from the SPAR model, we first draw an angle $\mathbf{w}$ from $f_{\mathbf{W}}$, then use inversion sampling to generate a corresponding value $z$ from the GP distribution with parameter vector $(\xi(\mathbf{w}),\sigma(\mathbf{w}))$, and finally define a radial value $r=u(\mathbf{w})+z$. The pair $(r,\mathbf{w})$ is then a random sample from the SPAR model. This can be converted back to the original variable space using the inverse transformation $\mathbf{x}=r\mathbf{w}$. As the SPAR model is only fitted to observations for which $r>u(\mathbf{w})$, one can create a sample (of the original random vector $\mathbf{X}$) of size $N$ by simulating $\zeta N$ points from the SPAR model, and then resampling $(1-\zeta)N$ points from observations with $r\leq u(\mathbf{w})$. The rationale for this is that there should be a sufficient number of observations within the body of the distribution to obtain a reasonable estimate from resampling.
2.2 Exceedance probability contours
As described in [mackay2025spar], the SPAR model provides an explicit means for calculating a contour with a specified exceedance probability $\beta\in(0,\zeta]$. The radius of the contour at angle $\mathbf{w}$ is simply the quantile of the GP distribution at exceedance probability $\beta/\zeta$, given by
$r_{\beta}(\mathbf{w})=u(\mathbf{w})+\frac{\sigma(\mathbf{w})}{\xi(\mathbf{w})}\left(\left(\frac{\beta}{\zeta}\right)^{-\xi(\mathbf{w})}-1\right).$ (9)
This contour is defined in terms of the probability of an observation falling anywhere outside the contour region, or the ‘total exceedance probability’. As such, these contours are more conservative than those defined in terms of marginal exceedance probabilities, such as IFORM contours (or variants thereof), with the conservatism increasing with the number of dimensions [mackay2021marginal]. Moreover, if the primary interest of the analysis is to estimate environmental contours, then the use of the SPAR model is not necessary. Instead, we recommend the use of the Direct-IFORM method [Derbanne2019, mackay2023-DIFORM], which does not require a model for the joint density or any assumptions about the dependence structure between the variables.
The contours defined in (9) can be projected into two dimensions. In Cartesian space, points on the contour are given by $\mathbf{x}=r_{\beta}(\mathbf{w})\,\mathbf{w}$ for $\mathbf{w}\in\mathbb{S}^{d-1}$. The radii $r_{\beta}\geq 0$ can be computed for a discrete set of points on the sphere (see Appendix A), and the projection into dimensions $i,j\in\{1,...,d\}$ is just the $i$ and $j$ components of $\mathbf{x}$.
However, in our application it was found that these contours do not provide a useful diagnostic tool. This is because they do not account for the angular density. In regions of low angular density, estimates of the contours can be highly uncertain. Simpson and Tawn [simpson2024inference] proposed to mitigate for this by either setting the radius to zero when the angular density is zero, or adjusting the exceedance level to account for the angular density. However, the latter approach requires an integral over angles that becomes impractical in higher dimensions. We have therefore opted to use alternative diagnostics, described in Section 4.5.
@@ 3 Inference
Inference for the SPAR model involves estimating the angular density, and the GP threshold and parameter functions. Since the angular-radial density factorises (see (5)), these problems are separable: inference for $f_{\mathbf{W}}$ can be conducted independently of that for $(u(\mathbf{w}),\sigma(\mathbf{w}),\xi(\mathbf{w}))$. Inference for the angular density is discussed in Section 3.1, and modelling of the conditional radial variable is discussed in Section 3.2. Code for fitting our model is available upon reasonable request.